OpenStudy (anonymous):

How many mph is a bullet from a pistol traveling if it leaves the muzzle at 1200 feet per second? The answer is 818 but I need to know the math behind the answer. please help thanks!

OpenStudy (anonymous):

The bullet is traveling \[\frac{1200 feet}{second}\] We need to turn this into miles instead of feet and hours instead of seconds. Let's convert it to hours first. There are 60 seconds in a minute and 60 minutes in a hour. 60*60 = 3600 so there are 3600 seconds in an hour. Now we have\[\frac{1200feet}{second}\times\frac{3600 seconds}{hour} = \frac{4320000 feet}{hour}\] Now convert feet to miles. There are 5280 feet in a mile, so now we have:\[\frac{4320000feet}{hour}\times\frac{mile}{5280feet}=\frac{4320000miles}{5280hour}\] \[\frac{4320000miles}{5280hours} = \frac{181.\overline{18} miles}{hour}\] or about 181 miles/hour.

OpenStudy (anonymous):

Whoops! I meant \[\frac{818.\overline{18} miles}{hour}\] or about 818 miles/hour Sorry about that